OpenStudy (anonymous):

integrate 1/ (t^3 (Square root (t^2-4) (dt) with a range of 4, (2 square root of 2). These ranges are next to the integrate sign. If anyone one could help me this would be great. Thanks.

OpenStudy (anonymous):

trig sub for this one

OpenStudy (anonymous):

it would be the sec trig sub correct. And if so my question is with the numbers well ranges that are given how do i work this into the problem.

OpenStudy (anonymous):

change the limits of integration when you make the sub then you don't have to change back

OpenStudy (anonymous):

Okay. I did that and got 4 to change into Pie over 3 but i am confused about the 2 square root of 2 how this one changes. I feel i have it wrong.

OpenStudy (anonymous):

sub is \(x=\sec(\theta)\) i think so \(\theta=\sec^{-1}(x)\)

OpenStudy (anonymous):

Oh okay so would that mean my pie over 3 is incorrect?

OpenStudy (anonymous):

\[\int _4^{2\sqrt{2}}\frac{dt}{t^3\sqrt{t^2-1}}\] is that it?

OpenStudy (anonymous):

Yes but the ranges are swapped.

OpenStudy (anonymous):

oh and the 1 is a 4 that is in the square root under the denominator

OpenStudy (anonymous):

or maybe this \[\int _{2\sqrt{2}}^4\frac{dt}{t^3\sqrt{t^2-1}}\]

OpenStudy (anonymous):

oooh ok that is better \[\int _{2\sqrt{2}}^4\frac{dt}{t^3\sqrt{t^2-4}}\]

OpenStudy (anonymous):

yes this would be my problem

OpenStudy (anonymous):

that makes the sub \[x=2\sec(\theta)\] and so \(\theta=\sec^{-1}(\frac{x}{2})\)

OpenStudy (anonymous):

that makes a lot more sense now you can compute the inverse easier

OpenStudy (anonymous):

thank you. This should now help me figure out where i have gone wrong within my problem. So thank you.

OpenStudy (anonymous):

lower limit is easy right? i get \(\frac{\pi}{4}\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yes i got \[\pi/4\] for the bottom

OpenStudy (anonymous):

k good top is probably easy too my guess \(\frac{\pi}{3}\) without even looking as the problem as been cooked up to be easy (sort of)

OpenStudy (anonymous):

Ya i got that too. But i am stuck at the ending if what i got is on the right track i dont know how to take it further to get the answer which i was given on my homework site. I feel like i am missing something, just dont know what?

OpenStudy (anonymous):

what did you get when you made the sub?

OpenStudy (anonymous):

\[\frac{2\sec(\theta)\tan(\theta)}{16\sec^3(\theta)\tan(\theta)}\] i think is the first step

OpenStudy (anonymous):

Yes i got this part

OpenStudy (anonymous):

leaving \[\frac{1}{8}\cos^2(\theta)\]

OpenStudy (anonymous):

Yes this is where i am now confused on how to take this further.

OpenStudy (anonymous):

oooh this is the easy gimmick

OpenStudy (anonymous):

\[\cos^2(\theta)=\frac{1}{2}(1+\cos(2\theta))\]

OpenStudy (anonymous):

and that is easy to integrate

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

okay sorry to bother you again. Is there anyway you can walk me through the rest of the steps for the problem cause i thought i saw were i needed to take this problem but it is not working out for me. If you could help i would appreciate it.

OpenStudy (anonymous):

Never mind i figured it out thanks though. And again thanks for all the help.